OpenStudy (anonymous):

how do you solve this? x over 5 is greater than or equal to -2

OpenStudy (anonymous):

\[\frac{x}{5}\geq -2\] \[x\geq-10\] done in one step

OpenStudy (anonymous):

how you get the -10 though ? what do you have to multiply or add